# Searching for light long-lived particles at SHiNESS

### v2

@@ 2.2 Electrophilic ALPs with weak-violating couplings
where $g$ is the SU(2) coupling constant, $c_{ee}$ is a dimensionless coupling between the ALP and the electrons, and $m_{e}$ is the electron mass. $\Lambda$ is the effective cut-off scale of new physics and $P_{L}$ is the left-chiral projector. The dots denote a set of terms arising from chiral anomaly and irrelevant to the phenomenology in our study, and are therefore ignored here.
@@
where $m_{a}$ is the ALP mass. Following the practice taken in Ref. Altmannshofer:2022ckw , we assume that the ALP decays with a 100% branching ratio into an electron-positron pair, and ignore the loop-induced photon-pair decay mode (which is sub-dominant for the relevant mass range anyway).
The existing bounds on $c_{ee}/\Lambda$ in the weak-violating scenario are obtained from various laboratory and astrophysical searches and observations BaBar:2014zli ; Konaka:1986cb ; Riordan:1987aw ; Bjorken:1988as ; Bross:1989mp ; Morel:2020dww ; Parker:2018vye ; Lucente:2021hbp ; KTeV:2003sls ; LHCb:2015ycz ; NA62:2020xlg ; SINDRUM:1986klz ; Poblaguev:2002ug ; CHARM:1985anb . In addition, rough potential projections of sensitivities, expected at the PIONEER experiment PIONEER:2022yag , kaon factories Goudzovski:2022vbt , as well as the LHC, are also available and reproduced from Ref. Altmannshofer:2022ckw . In particular, the latter corresponds to a benchmark value of $10^{-5}$ for the ratio $\text{Br}_{W}=\text{Br}(W^{+}\to l^{+}\nu_{l}a)/\text{Br}(W^{+}\to e^{+}\nu)$, expected at the future LHC.
At the end, we comment that Ref. Altmannshofer:2022ckw also studied a weak-preserving scenario where the contributions from the four-point vertex are vanishing. We explicitly checked and found that in this case the expected sensitivity reach of SHiNESS is already all excluded by the SLAC E137 experiment Bjorken:1988as , if we restrict ourselves to the ALP coupled to electrons only, as the strong helicity suppression largely reduces the productions rates of the ALP. If the ALP is coupled to muons only, it will be produced in $\pi^{+}\to\mu^{+}\nu_{\mu}a$ decays; in this case, the helicity suppression is negligible, but SHiNESS will still have no sensitivity, for kinematic reasons related to cut selections to be imposed on final-state leptons (see discussion in Sec. 3). Therefore, we do not include this weak-preserving scenario in the present study.
@@ 3 Experimental setup
Simulation indicates a neutrino yield of 0.3 per flavor ($\nu_{\mu},\nu_{e},$ and $\bar{\nu}_{\mu}$) per proton, implying $8.5\times 10^{22}$ neutrinos per flavor per year produced from $\pi^{+}$ decays at rest Baxter:2019mcx . Since the $\pi^{+}$ mesons decay to $\mu^{+}\nu_{\mu}$ with a branching fraction of $\sim 99.99\%$ ParticleDataGroup:2022pth , this corresponds to $8.5\times 10^{22}$ $\pi^{+}$ mesons produced each year and decaying at rest. The $\mu^{+}$ leptons produced from $\pi^{+}$ decays also undergo ionization processes and quickly decay at rest in the close vicinity Baxter:2019mcx ; Burman:2003ek .
@@
The proposal of the SHiNESS detector focuses on the $e^{-}e^{+}$-pair signature in the context of searches for LLPs such as long-lived HNLs. This aim is realized by installing arrays of photomultipliers. Further, with the ability to detect both Cherenkov photons and scintillation light, the SHiNESS detector has strong vertex resolution and can reconstruct particle directions. This is further enabled by distinguishing between the Cherenkov and scintillation light that is allowed by various strategies and hardware implementation such as photosensors for their excellent timing resolution. We refer to Ref. Soleti:2023hlr for more detail. In this work, as already specified in the previous section, we also concentrate on the signature of excesses in the number of $e^{-}e^{+}$ events (coming from decays of long-lived binos or ALPs).
@@
We follow Ref. Soleti:2023hlr for a brief discussion on background estimates. The LLP search requires reconstruction of a vertex formed from two charged leptons. The two lepton signals should be separated in time within 100 ns, and both should have an energy larger than 17 MeV. In addition, the opening angle between the two leptons should be at least $15^{\circ}$. With these requirements, the main background sources are cosmic ray, charged-current $\nu_{e}$ interactions, neutral-current $\nu$ interactions, and $\bar{\nu}_{e}$ beam component. These sources sum up to about 61 background events per year. A large, $50\%$ systematic uncertainty is assigned, corresponding to 31 events per year Soleti:2023hlr , and we take the same assumption in this work.
For this study, we take 2 years’ operation time as a benchmark, and calculate the signal-event number for the exclusion bounds at 90% confidence level (C.L.) below. For two years’ data-collection time, we have $122\pm 43.84$ background events in total. Performing an unbinned $\chi^{2}$ test at 90% C.L., we find the corresponding bound on NP signal-event number to be $N_{S}^{90}\approx 72.17$. For detail of this estimate, we refer to Appendix A.
We perform simulations in order to estimate the detector acceptance to the long-lived binos or ALPs, as well as the cut efficiencies on the energies and opening angle of the final-state electron and positron. Specifically, the kinematics of the final-state particles in two-body decays can be simply determined by the masses of the mother and daughter particles. For three-body decays, kinematically allowed, random values of transferred momentum squared are taken in order to compute the 4-momenta of the final-state particles. Additionally, the particles are boosted to the laboratory frame. These procedures allow to simulate the kinematics of the LLPs and their decay products including the final-state electron and positron. With the knowledge of these particles’ kinematics, we proceed to compute the detector acceptance and cut efficiencies.
@@
For each simulated event, we check first on the final-state electron and positron whether they pass the above-mentioned cuts. If the energy and opening-angle requirements are satisfied, the associated simulated LLP is taken for computing its decay probability inside the active volume of the SHiNESS detector. The probability of the $i^{\text{th}}$ simulated LLP to pass the lepton cuts and to decay inside the fiducial volume of SHiNESS is calculated as follows
@@
where we include the cut requirements on the final-state electron and positron, and “f.v.” in $\epsilon^{\text{f.v.}}_{i^{\text{th}}\text{ LLP}}$ stands for “fiducial volume”, and “acc.” is abbreviation for “acceptance”. To compute $\epsilon^{\text{f.v.}}_{i^{\text{th}}\text{ LLP}}$, we assume that in the vertical dimension the target is located in the middle of the SHiNESS detector.
@@
In Fig. 1 we display a bird’s-eye view of the experimental setup, where $\theta=35^{\circ}$ and $D=25$ m is the distance of the detector to the tungsten target. The center of SHiNESS is at $y=0$. We consider only the positive-$z$ hemisphere (requiring the $z$-component of the momentum $p_{z}>0$ on the LLPs) and use the following formulas to compute $\epsilon^{\text{f.v.}}_{i^{\text{th}}\text{ LLP}}$,
$\epsilon^{\text{f.v.}}_{i^{\text{th}}\text{ LLP}}$ $=$ $\displaystyle\sum_{j=1}^{2}\delta\phi^{j}\cdot e^{-L^{j}_{i}/\lambda_{i}^{z}}\cdot\Big{(}1-e^{-\frac{L^{\prime j}_{i}}{\lambda_{i}^{z}}}\Big{)}\,,$ (14)
$L^{j}_{i}$ $=$ $\text{min}\bigg{(}\text{max}\bigg{(}d_{h},\frac{d_{v}^{j}}{\tan{\theta_{i}}}\bigg{)},d_{h}+L_{h}\bigg{)}\,,$ (15)
$L^{\prime j}_{i}$ $=$ min($\text{min}\bigg{(}\text{max}\bigg{(}d_{h},\frac{d_{v}^{j}+R}{\tan{\theta_{i}}}\bigg{)},d_{h}+L_{h}\bigg{)}-L^{j}_{i},\,\,\,\,\,\,\,\,\,\,\,\,\,\,$ (16)
$\delta\phi^{1}$ $=$ $2\cdot\arctan{\Big{(}\frac{L/2}{R/2+d_{v}^{1}}\Big{)}}/(2\pi),$ (17)
$\delta\phi^{2}$ $=$ $2\cdot\arctan{\Big{(}\frac{L/2}{R/2+d_{v}^{2}}\Big{)}}/(2\pi),$ (18)
$d_{v}^{1}$ $=$ $D\cdot\sin{35^{\circ}}-R,$ (19)
$d_{v}^{2}$ $=$ $D\cdot\sin{35^{\circ}},$ (20)
where $L=4$ m and $R=2$ m are the height and radius of the fiducial volume, respectively. $d_{h}=D\cdot\cos{35^{\circ}}$ is the longitudinal (along the $z$-axis) distance from the fixed target to the SHiNESS detector, and $d_{v}^{1}$ and $d_{v}^{2}$ are the distances along the $x$-axis from the target to the near end and middle position of SHiNESS, respectively. $\lambda_{i}^{z}=\beta_{i}^{z}\gamma_{i}c\tau$ is the boosted decay length of the $i^{\text{th}}$ simulated LLP in the $z$-direction with $\beta_{i}^{z}$ being the absolute speed, $\gamma_{i}$ the boost factor, and $c\tau$ the proper decay length. $\delta\phi^{j}$ is the detector coverage in the azimuthal direction. The general idea of Eqs. (14) is to divide the detector into two parts (corresponding to $j=1,2$), one $(j=2)$ above the other $(j=1)$ along the $x$-axis, compute the LLP’s decay probability inside each part, and sum them up in the end. While Eq. (14) is only an approximate formula, this approach of considering two parts reduces the error.
We note that despite the target wheel being 2.6-m wide Garoby:2017vew , we have assumed above, for simplicity, that the target is point-like. Moreover, both $\pi^{+}$ mesons and $\mu^{+}$ leptons are assumed to decay at rest at the interaction point on the target, considering the ionization effects as mentioned above. We expect this simplification of the setup to affect the final results to a rather minor extent (see discussion in Ref. Baxter:2019mcx ).
@@
where $N_{\text{LLP}}$ is the total number of the LLPs expected to be produced in two years’ operation of the ESS, $\text{Br}(\text{LLP}\to e^{-}e^{+}X)$ is the decay branching ratio of the LLP into an electron-positron pair plus anything, and $\epsilon^{\text{cut\& acc.}}_{\text{avg.}}$ is evaluated with the following expression:
@@ 4 Numerical results
Following the discussion in the previous section, we present numerical results in this section, for the exclusion bounds at 90% C.L.
@@ 4.1 Light binos
For the theoretical model of long-lived light neutralinos in the RPV-SUSY, despite the multiple benchmark scenarios relevant in principle for LLP searches at SHiNESS as listed in Table 1, SHiNESS can probe large, unexcluded parameter space only in some of them.
@@
For benchmarks B1.1 and B1.2, strong bounds exist on the single coupling $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$ as functions of the light bino’s mass, from $0\nu\beta\beta$ decays and HNL searches, cf. the discussion in Sec. 2. Additionally, for B1.1, the product bound given in Eq. (8) constrains the scenario tightly. Therefore, we choose to present only a sensitivity plot in Fig. 2 for B1.2, shown in the plane $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$ vs. $\lambda_{131}/m^{2}_{\tilde{f}}$, where we fix the light bino’s mass at 20 MeV (black), 70 MeV (scarlet), and 120 MeV (green). The horizontal dashed curves are current upper bounds on $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$, where we have picked, for each bino-mass choice, the stronger limit between that derived from $0\nu\beta\beta$ decays and that from HNL searches. The vertical red curves are present upper limits on $\lambda_{131}/m^{2}_{\tilde{f}}$ for $m_{\tilde{f}}=1$ TeV and 2 TeV (see Eq. (5)) which do not depend on $m_{\tilde{\chi}^{0}_{1}}$. We find that for all three choices of the bino mass, a triangle-shaped window that is still allowed in the shown parameter space can be excluded by SHiNESS in two years.
@@
Figure 2: Sensitivity reach of SHiNESS in two years’ data-collection time to the benchmark scenario B1.2, shown in the $(\lambda_{131}/m^{2}_{\tilde{f}},\lambda^{\prime}_{111}/m^{2}_{\tilde{f}})$ plane, for $m_{\tilde{\chi}^{0}_{1}}=20$ MeV, 70 MeV, and 120 MeV. The horizontal dashed curves are present bounds on $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$ obtained from $0\nu\beta\beta$ decays Mohapatra:1986su ; Hirsch:1995zi ; Hirsch:1995ek ; Babu:1995vh and recasting HNL searches Dreiner:2023gir , and the vertical dashed red curve are current limits on $\lambda_{131}/m^{2}_{\tilde{f}}$ (see Eq. (5)).
The benchmark scenarios B2.1 and B2.2 both assume a coupling $\lambda^{\prime}_{211}$ leading to the process $\pi^{+}\to\mu^{+}\tilde{\chi}^{0}_{1}$ producing the light bino. In this case, the light bino is produced with the maximally possible energy $m_{\pi^{+}}-m_{\mu^{+}}\approx 35$ MeV, since the $\pi^{+}$’s decay at rest. As a result, it is impossible for the electron-positron pair produced in the decays of the light bino to pass the lepton-cut selections required in the search, and SHiNESS hence is insensitive to these benchmarks. Note that, of course, since in B2.1 the coupling $\lambda_{121}$ is present which can induce both $\mu^{+}\to\tilde{\chi}^{0}_{1}\bar{\nu}_{e}e^{+}$ and $\tilde{\chi}^{0}_{1}\to\overset{(-)}{\nu}_{\mu}e^{+}e^{-}$ decays, SHiNESS should have sensitivities, but this decay chain reduces the scenario to B3.1 and is hence not taken into account here.
The benchmark B3.1 is interesting in that it is the only single-coupling scenario listed in Table 1, where the only non-vanishing coupling $\lambda_{121}$ mediates both the production and decay of the lightest neutralino. We show a sensitivity plot in Fig. 3 in the $(m_{\tilde{\chi}^{0}_{1}},\lambda_{121}/m^{2}_{\tilde{f}})$ plane.
Figure 3: Sensitivity reach of SHiNESS to the benchmark B3.1, displayed in the plane $\lambda_{121}/m^{2}_{\tilde{f}}$ vs. $m_{\tilde{\chi}^{0}_{1}}$. The red dashed horizontal curves are current limits on $\lambda_{121}/m^{2}_{\tilde{f}}$, cf. Eq. (4), for $m_{\tilde{f}}=1$ TeV and 2 TeV, and the gray dashed curve is the bound extracted from Ref. Dreiner:2023gir .
@@
Here, SHiNESS in two year’s operation can probe $m_{\tilde{\chi}^{0}_{1}}$ up to right below the muon threshold. Compared to the existing bounds from charged-current universality (red horizontal curves), cf. Eq. (4), and those from the uncertainty on the measured decay width of $\mu^{\pm}\to e^{\pm}+\text{invisible}$ Dreiner:2023gir (gray curve), we find that SHiNESS can exclude large parameter space for $m_{\tilde{\chi}^{0}_{1}}$ roughly between 20 MeV and 100 MeV, probing $\lambda_{121}/m^{2}_{\tilde{f}}$ down to as low as about $1.5\times 10^{-8}$ GeV-2.
@@
B3.2 is the only benchmark (besides B3.1) where the production coupling also induces kinematically-allowed bino decays.
Figure 4: Sensitivity reach of SHiNESS to the benchmark B3.2. “C.C. universality” stands for “charged-current universality”, cf. Eq. (4).
@@
We choose to present the numerical results in Fig. 4, with two plots shown in the plane $\lambda_{121}/m^{2}_{\tilde{f}}$ vs. $\lambda_{131}/m^{2}_{\tilde{f}}$ and $\lambda_{121}/m^{2}_{\tilde{f}}=\lambda_{131}/m^{2}_{\tilde{f}}$ vs. $m_{\tilde{\chi}^{0}_{1}}$, respectively. The present limits on both $\lambda_{121}/m^{2}_{\tilde{f}}$ (Eq. (4)) and $\lambda_{131}/m^{2}_{\tilde{f}}$ (Eq. (5)) are overlaid with blue and red curves, respectively, for both sfermion masses of 1 TeV and 2 TeV. Moreover, for $\lambda_{121}/m^{2}_{\tilde{f}}$, the bounds obtained by considering $\sigma_{\Gamma(\mu^{\pm}\to e^{\pm}+\text{invis.})}$ are also shown, in dashed lines with different colors corresponding to the chosen mass values (10 MeV, 20 MeV, and 60 MeV) of the light bino (upper panel) and in a gray solid line (lower panel). In the upper plot of Fig. 4, we observe that for small values of $\lambda_{131}/m^{2}_{\tilde{f}}$ the sensitivity curves become flat, reflecting the fact that the $\lambda_{121}$ coupling alone can lead to sufficiently many signal events. We observe that for $m_{\tilde{\chi}^{0}_{1}}=20$ MeV, 60 MeV, and 100 MeV, SHiNESS can probe large parameter space beyond the present bounds. Further, for increasing $m_{\tilde{\chi}^{0}_{1}}$ the sensitivity of SHiNESS is enhanced, except for the $m_{\tilde{\chi}^{0}_{1}}=100$ MeV case where the phase-space suppression effect comes into play. The lower plot of Fig. 4, similar to Fig. 3, shows that SHiNESS can extend the current bounds largely in just two years, mainly for $m_{\tilde{\chi}^{0}_{1}}$ between about 20 MeV and 100 MeV.
For B4.1, the product bound as given in Eq. (7) is so strong that it has excluded all the parameter space that would be probed by SHiNESS if we set $\lambda_{121}=\lambda_{122}$. If the two non-vanishing $\lambda$ couplings are independent, when $\lambda_{121}/m^{2}_{\tilde{f}}$ is sufficiently small, SHiNESS can only probe $\lambda_{122}/m^{2}_{\tilde{f}}$ values that are already excluded by the product bound. For larger values of $\lambda_{121}/m^{2}_{\tilde{f}}$, the search is insensitive to $\lambda_{122}$ as the single coupling $\lambda_{121}$ leading to both production and decay of the light bino alone would result in sufficiently many signal events (see Fig. 3). For these reasons, we do not present numerical results for B4.1.
@@
The results of benchmarks B5.1-B6.1 are all similar to those of the previous benchmarks, and we therefore refrain from showing them. We decide to present numerical results for B6.2, where there is no available bound from reinterpretation of the HNL searches or $\sigma_{\Gamma(\mu^{\pm}\to e^{\pm}+\text{invis.})}$. We note that we choose not to show the results of B4.2 because their conclusions are close to those of B6.2.
@@
Figure 5: Sensitivity reach of SHiNESS to the benchmark B6.2, shown in the plane $\lambda_{312}/m^{2}_{\tilde{f}}$ vs. $\lambda_{131}/m^{2}_{\tilde{f}}$ (upper plot) and $\lambda_{312}/m^{2}_{\tilde{f}}=\lambda_{131}/m^{2}_{\tilde{f}}$ vs. $m_{\tilde{\chi}^{0}_{1}}$ (lower plot).

### v3

@@ 2.2 Electrophilic ALPs with weak-violating couplings
where $g$ is the SU(2) coupling constant, $c_{ee}$ is a dimensionless coupling, and $m_{e}$ is the electron mass. $\Lambda$ is the effective cut-off scale of new physics and $P_{L}$ is the left-chiral projector. The dots denote a set of terms arising from chiral anomaly and irrelevant to the phenomenology in our study, and are therefore ignored here.
@@
where $m_{a}$ is the ALP mass. Following the practice taken in Ref. Altmannshofer:2022ckw , we assume that the ALP decays with a 100% branching ratio into an electron-positron pair, and ignore the loop-induced photon-pair decay mode (which is sub-dominant for the relevant mass range anyway). It is worth noting that both the $\pi^{+}\to e^{+}\,\nu_{e}\,a$ and $a\to e^{+}e^{-}$ decays are mediated by the $c_{ee}$ coupling.
The existing bounds on $c_{ee}/\Lambda$ in the weak-violating scenario are obtained from various laboratory and astrophysical searches and observations BaBar:2014zli ; Konaka:1986cb ; Riordan:1987aw ; Bjorken:1988as ; Bross:1989mp ; Morel:2020dww ; Parker:2018vye ; Lucente:2021hbp ; KTeV:2003sls ; LHCb:2015ycz ; NA62:2020xlg ; SINDRUM:1986klz ; Poblaguev:2002ug ; CHARM:1985anb . In addition, rough potential projections of sensitivities, expected at the PIONEER experiment PIONEER:2022yag , kaon factories Goudzovski:2022vbt , as well as the LHC, are provided in Ref. Altmannshofer:2022ckw . In particular, the latter corresponds to a benchmark value of $10^{-5}$ for the ratio $\text{Br}_{W}=\text{Br}(W^{+}\to l^{+}\nu_{l}a)/\text{Br}(W^{+}\to e^{+}\nu)$, expected at the future LHC.
At the end, we comment that Ref. Altmannshofer:2022ckw also studied a weak-preserving scenario where the contributions from the four-point vertex are vanishing. We explicitly checked and found that in this case the expected sensitivity reach of SHiNESS is already all excluded by the SLAC E137 experiment Bjorken:1988as , if we restrict ourselves to the ALP coupled to electrons only, as the strong helicity suppression largely reduces the productions rates of the ALP. If the ALP is coupled to muons only, it will be produced in $\pi^{+}\to\mu^{+}\nu_{\mu}a$ decays; in this case, the helicity suppression is negligible, but SHiNESS will still have no sensitivity, for kinematic reasons related to cut selections to be imposed on final-state leptons (see discussion in Sec. 3). Therefore, we do not include weak-preserving scenarios in the present study.
@@ 3 Experimental setup
Simulation predicts an expected annual rate of $8.5\times 10^{22}$ decays at rest for both $\pi^{+}$ mesons and $\mu^{+}$ leptons at the ESS Baxter:2019mcx .
@@
The proposal of the SHiNESS detector focuses on the $e^{-}e^{+}$-pair signature in the context of searches for LLPs such as long-lived HNLs. This aim is realized by installing arrays of photomultipliers. Further, with the ability to detect and distinguish between both Cherenkov photons and scintillation light, the SHiNESS detector has strong vertex resolution and can reconstruct particle directions. In particular, the separation of the Cherenkov and scintillation light is allowed by various strategies and hardware implementation such as photosensors for their excellent timing resolution. Thus, the electron and the positron in a signal event can be separated, and the energy resolution of the charged leptons is at the $\mathcal{O}(1)\%$ level, cf. Fig. 8 of Ref. Soleti:2023hlr . We refer to Ref. Soleti:2023hlr for more detail. In this work, as already specified in the previous section, we also concentrate on the signature of excesses in the number of $e^{-}e^{+}$ events (coming from decays of long-lived binos or ALPs).
@@
We follow Ref. Soleti:2023hlr for a brief discussion on background estimates. The LLP search requires reconstruction of a vertex formed from two charged leptons. For this purpose, the two lepton signals should be separated in time within 100 ns, and both should have an energy larger than 17 MeV. In addition, the opening angle between the two leptons should be at least $15^{\circ}$  [^4] The lepton cuts imply that the light bino produced in $\pi^{+}\to\mu^{+}\tilde{\chi}^{0}_{1}$ decays at rest via the coupling $\lambda^{\prime}_{211}$ cannot be probed by SHiNESS because the light bino would have at most an energy of $m_{\pi^{+}}-m_{\mu^{+}}\approx 35$ MeV. As a result, it is impossible for the electron-positron pair produced in the decays of the light bino to pass the lepton-cut selections required in the search. For the same reason, SHiNESS is insensitive to the ALP produced in $\pi^{+}\to\mu^{+}\,\nu_{\mu}\,a$ decays.. With these requirements, the main background sources are cosmic ray, charged-current $\nu_{e}$ interactions, neutral-current $\nu$ interactions, and $\bar{\nu}_{e}$ beam component. These sources sum up to about 61 background events per year. A large, $50\%$ systematic uncertainty is assigned, corresponding to 31 events per year Soleti:2023hlr , and we take the same assumption in this work. For a benchmark 2-year operation time, we thus expect $122\pm 43.84$ background events in total, corresponding to an NP signal-event number of 72.17 for exclusion bounds at 90% confidence level (CL).
@@
Using Monte-Carlo techniques, we estimate the efficiency of the LLP decay products fulfilling a specified list of criteria. For each simulated event, we check first on the final-state electron and positron whether they pass the above-mentioned cuts. If the energy and opening-angle requirements are satisfied, the associated simulated LLP is taken for computing its decay probability inside the active volume of the SHiNESS detector. The probability of the $i^{\text{th}}$ simulated LLP to pass the lepton cuts and to decay inside the fiducial volume of SHiNESS is calculated as follows
@@
where we include the cut requirements on the final-state electron and positron, “f.v.” in $\epsilon^{\text{f.v.}}_{i^{\text{th}}\text{ LLP}}$ stands for “fiducial volume”, and “acc.” is abbreviation for “acceptance”. To compute $\epsilon^{\text{f.v.}}_{i^{\text{th}}\text{ LLP}}$, we assume that the target is point-like and located in the middle of the SHiNESS detector in the vertical dimension.
@@
In Fig. 1 we display a bird’s-eye view of the experimental setup, where $\theta=35^{\circ}$ and $D=25$ m is the distance of the detector to the tungsten target. The center of SHiNESS is at $y=0$. With the Monte-Carlo simulations, we calculate the decay-in-volume probability, $\epsilon^{\text{f.v.}}_{i^{\text{th}}\text{ LLP}}$, by taking into account the kinematics and the lifetime of the $i^{\text{th}}$ simulated LLP, as well as the previously defined fiducial volume, if it has satisfied the lepton-cut requirements.
@@
where $N_{\text{LLP}}$ is the total number of the LLPs expected to be produced in two years’ operation of the ESS, $\text{Br}(\text{LLP}\to e^{-}e^{+}X)$ is the decay branching ratio of the LLP into an electron-positron pair plus anything, and the average efficiency $\epsilon^{\text{cut\& acc.}}_{\text{avg.}}$ is evaluated with the following expression:
@@
Figure 2: Lepton-cut efficiencies as functions of the mass of the bino or ALP, where only the requirements on the energies of the final-state electron and positron, as well as the opening angle between them, are included. The blue (scarlet) curve is for the bino benchmark B1 (B2-B4), and the green one is for the electrophilic ALP model.
In Fig. 2, we show a plot for the average efficiency of the lepton cuts as functions of the LLP mass, where the fiducial-volume requirement is not included. The blue and scarlet curves correspond to the bino benchmarks B1 and B2-B4, respectively, and the green one is for the electrophilic ALP scenario. The latter scenario has the strongest lepton-cut efficiencies, mainly because the ALP undergoes two-body decays into an electron-positron pair, while the light binos decay into three particles among which the active neutrino carries away energy resulting in less energetic electron and positron. We find that the lepton-cut efficiencies lie in $\mathcal{O}(10)\%$ for almost the whole range of the kinematically allowed LLP mass. However, the total efficiency $\epsilon^{\text{cut\&acc.}}_{\text{avg.}}$ depends also on the lifetime of the LLP since it includes the fiducial-volume requirement. In our numerical simulation, we observe that the maximal value of $\epsilon^{\text{cut\&acc.}}_{\text{avg.}}$ is about $10^{-4}$ in all theoretical scenarios.
@@ 4 Numerical results
We present numerical results in this section, for the exclusion bounds at 90% CL.
@@ 4.1 Light binos
For the theoretical model of long-lived light neutralinos in the RPV-SUSY, SHiNESS can probe large, unexcluded parameter space in the benchmark scenarios listed in Table 1.
@@
For benchmark B1, strong bounds exist on the single coupling $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$ as functions of the light bino’s mass, from $0\nu\beta\beta$ decays and HNL searches, cf. the discussion in Sec. 2. We present a sensitivity plot for B1 in the left panel of Fig. 3, shown in the $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$ vs. $\lambda_{131}/m^{2}_{\tilde{f}}$ plane, where we fix the light bino’s mass at 20 MeV (blue), 70 MeV (scarlet), and 120 MeV (green). The horizontal dashed curves are current upper bounds on $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$, where we have picked, for each bino-mass choice, the stronger limit between that derived from $0\nu\beta\beta$ decays and that from HNL searches. The vertical solid red curves are present upper limits on $\lambda_{131}/m^{2}_{\tilde{f}}$ for $m_{\tilde{f}}=1$ TeV and 2 TeV derived from measurements on $R_{\tau}$ (see Table 2) which do not depend on $m_{\tilde{\chi}^{0}_{1}}$. These present limits are labeled with the processes from which they were obtained. We find that for all the three choices of the bino mass, a triangle-shaped window that is allowed in the shown parameter space can be excluded by SHiNESS in two years.
@@
Figure 3: Left: sensitivity reach of SHiNESS at 90% CL in two years’ data-collection time to the benchmark scenario B1, shown in the $(\lambda_{131}/m^{2}_{\tilde{f}},\lambda^{\prime}_{111}/m^{2}_{\tilde{f}})$ plane, for $m_{\tilde{\chi}^{0}_{1}}=20$ MeV, 70 MeV, and 120 MeV. The horizontal dashed curves are present bounds on $\lambda^{\prime}_{111}/m^{2}_{\tilde{f}}$ obtained from $0\nu\beta\beta$ decays Mohapatra:1986su ; Hirsch:1995zi ; Hirsch:1995ek ; Babu:1995vh and recasting HNL searches Dreiner:2023gir with the excluded regions shaded with the colors of the corresponding masses of $\tilde{\chi}^{0}_{1}$, and the vertical solid red curves are current limits on $\lambda_{131}/m^{2}_{\tilde{f}}$ for sfermion masses of 1 TeV and 2 TeV obtained by measuring $R_{\tau}$, with the area to the right of these vertical lines being already excluded, cf. Table 2. Right: sensitivity reach of SHiNESS to the benchmark B2, displayed in the $\lambda_{121}/m^{2}_{\tilde{f}}$ vs. $m_{\tilde{\chi}^{0}_{1}}$ plane. The red solid horizontal curves are current limits on $\lambda_{121}/m^{2}_{\tilde{f}}$ obtained by considering C. C. universality, cf. Table 2, for $m_{\tilde{f}}=1$ TeV and 2 TeV, and the third red solid curve is the bound obtained by considering $\sigma_{\Gamma(\mu^{\pm}\to e^{\pm}+\text{invis.})}$.
@@
The benchmark B2 is the only single-coupling scenario listed in Table 1, where the only non-vanishing coupling $\lambda_{121}$ mediates both the production and decay of the lightest neutralino. We show a sensitivity plot in the right panel of Fig. 3 in the $(m_{\tilde{\chi}^{0}_{1}},\lambda_{121}/m^{2}_{\tilde{f}})$ plane. Here, SHiNESS in two year’s operation can probe $m_{\tilde{\chi}^{0}_{1}}$ up to right below the muon threshold. Compared to the existing bounds from C. C. universality Barger:1989rk (red horizontal curves), and those from the uncertainty on the measured decay width of $\mu^{\pm}\to e^{\pm}+\text{invisible}$ Dreiner:2023gir (the third red curve), we find that SHiNESS can exclude large parameter space for $m_{\tilde{\chi}^{0}_{1}}$ roughly between 20 MeV and 100 MeV, probing $\lambda_{121}/m^{2}_{\tilde{f}}$ down to as low as about $1.5\times 10^{-8}$ GeV-2.
@@
B3 is the only benchmark (besides B2) where the production coupling also induces kinematically allowed bino decays.
Figure 4: Sensitivity reach of SHiNESS to the benchmark B3 presented in the $\lambda_{121}/m^{2}_{\tilde{f}}$ vs. $\lambda_{131}/m^{2}_{\tilde{f}}$ plane (left plot) and the $\lambda_{121}/m^{2}_{\tilde{f}}=\lambda_{131}/m^{2}_{\tilde{f}}$ vs. $m_{\tilde{\chi}^{0}_{1}}$ plane (right plot). In the left panel, the bounds from $\sigma_{\Gamma(\mu^{\pm}\to e^{\pm}+\text{invis.})}$ on $\lambda_{121}/m^{2}_{\tilde{f}}$ for $m_{\tilde{\chi}^{0}_{1}}=100$ MeV are unavailable in Ref. Dreiner:2023gir .
@@
We choose to present the numerical results in Fig. 4, with two plots shown in the $\lambda_{121}/m^{2}_{\tilde{f}}$ vs. $\lambda_{131}/m^{2}_{\tilde{f}}$ and $\lambda_{121}/m^{2}_{\tilde{f}}=\lambda_{131}/m^{2}_{\tilde{f}}$ vs. $m_{\tilde{\chi}^{0}_{1}}$ planes, respectively. The present limits on both $\lambda_{121}/m^{2}_{\tilde{f}}$ and $\lambda_{131}/m^{2}_{\tilde{f}}$, derived from C. C. universality and $R_{\tau}$, are overlaid with blue and red curves, respectively, for both sfermion masses of 1 TeV and 2 TeV. Moreover, for $\lambda_{121}/m^{2}_{\tilde{f}}$, the bounds obtained by considering $\sigma_{\Gamma(\mu^{\pm}\to e^{\pm}+\text{invis.})}$ are also shown, in dashed horizontal lines with different colors corresponding to the chosen mass values (10 MeV, 20 MeV, and 60 MeV) of the light bino (left panel) and in a blue solid line (right panel). In the left plot of Fig. 4, we observe that for small values of $\lambda_{131}/m^{2}_{\tilde{f}}$ the sensitivity curves become flat, reflecting the fact that the $\lambda_{121}$ coupling alone can lead to sufficiently many signal events. We find that for $m_{\tilde{\chi}^{0}_{1}}=20$ MeV, 60 MeV, and 100 MeV, SHiNESS can probe large parameter space beyond the present bounds. Further, for increasing $m_{\tilde{\chi}^{0}_{1}}$ the sensitivity of SHiNESS is enhanced, except for the $m_{\tilde{\chi}^{0}_{1}}=100$ MeV case where the phase-space suppression effect comes into play. The right plot of Fig. 4, similar to the right plot of Fig. 3, shows that SHiNESS can extend the current bounds largely in just two years, mainly for $m_{\tilde{\chi}^{0}_{1}}$ between about 20 MeV and 100 MeV.
@@
We present numerical results of B4 next, where there is no available bound from reinterpretation of the HNL searches or $\sigma_{\Gamma(\mu^{\pm}\to e^{\pm}+\text{invis.})}$.
@@
Figure 5: Sensitivity reach of SHiNESS to the benchmark B4, shown in the $\lambda_{312}/m^{2}_{\tilde{f}}$ vs. $\lambda_{131}/m^{2}_{\tilde{f}}$ plane (left plot) and the $\lambda_{312}/m^{2}_{\tilde{f}}=\lambda_{131}/m^{2}_{\tilde{f}}$ vs. $m_{\tilde{\chi}^{0}_{1}}$ plane (right plot).